Grade 12th passMechanics

A star rotates with a period of 30 days about an axis through its centre. After the star undergoes a supernova explosion, the stellar core, which had a radius 10000 km,collapses into a neutron star of radius 3.0 km. Determine the period of rotation of the neutron star.

Ashim Joshi
7 Years agoGrade 12th pass

1 Answer

Khimraj
7 Years ago
conservation of angular momentum J*ω .. since J = kr^2 , then : 
J1*ω1 = J2*ω2
k*r12*ω1 = k*r22*ω2 
r12*ω1  = r22*ω2
r1 = 104 km
r2 = 3 km
108*ω1 = 9*ω2
ω2/ω1 =  108/9
T1/T2 = 108/9
T2 = T1*9/108 days = 30*24*3600*9/108 sec = 0.2333 sec
Hope it clears..............